Write the equation of the line parallel to 3x - y = 2 and passing through (2, -3).
Answer by stanbon(75887)

Write the equation of the line parallel to 3x-y=2 and passing through (2,-3).
-------------------
Given line: y = 3x-2, so it has a slope of 3
-----------
ANY line parallel to it must have the same slope: m=3
------------
Equation Form: y = mx + b
You know y= -3 when x = 2, and you know m=3; solve for b.
-3 = 3*2 + b
b = -9
----------
EQUATION:
y = 3x - 9
==============
Cheers,
Stan H.